OpenStudy (anonymous):

A police car leaves in pursuit of a hold-up car 30 min after the latter has left scene of the crime a 75 mph. a. how fast must the police car go if it is to catch up the hold-up car in one hour? b. how far will the cars have traveled when they meet?

OpenStudy (anonymous):

a. first, figure out the distance the hold-up car has traveled. s=vt, v=75 mph, t =0.5 +1=1.5hrs. then, use v=s/t to figure out the speed of the police car.